OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6. y = 1/3 x + 14/3 y = 3x + 10 y = -3x - 2 y = - 1/3 x + 10/3

OpenStudy (yttrium):

hint: 1. slopes of parallel lines are equal. 2. perform the slope-intercept form of the line

OpenStudy (anonymous):

I still have no clue man

OpenStudy (yttrium):

do you know how to find the slope of a line>?

OpenStudy (anonymous):

The important part of the question is "is parallel to y = -3x + 6" the slope is mx in y=mx+b y=mx+b is the same as y=-3x+6 meaning that -3x is your slope When to functions are parallel, it simply means that both slope are the same. -3x in the problem looks like -3x in one of the answer options. There are none other that look like it so that means that the correct answer is y = -3x - 2

OpenStudy (anonymous):

Thank you guys